You will make the following investments for a trip around the world: $2,600 today, $4,000 at the end of year two, and $1500 at the end of year five. (A) How much will you have in six years if you can earn 4.2% on your investments? (B) What equivalent amount could you put away today as a lump sum and have the same amount in six years?
Answer #1
Future value = present value*(1+ rate)^time
Future value = 2600*(1+0.042)^6
Future value = 3327.97
Future value = present value*(1+ rate)^time
Future value = 4000*(1+0.042)^4
Future value = 4715.53
Future value = present value*(1+ rate)^time
Future value = 1500*(1+0.042)^1
Future value = 1563

Total future value = 3327.97+4715.53+1563=9606.5

Future value = present value*(1+ rate)^time
9606.5 = Present value*(1+0.042)^6
Present value = 7505.14